Cambridge Company purchased a truck on January 1, 2018. Cambridge paid $23,000 for the truck. The truck is expected to have a $3,500 residual value and a 5-year life. Cambridge has a December 31 fiscal year end. Using the double-declining balance method, how much is the 2019 depreciation expense? (Enter only whole dollar values.) Hint: what is the year 2 depreciation amount?
Answer #1

Depreciation for 2019= $ 5,520

Working

Double declining Method
Cost $ 23,000
B Residual Value $ 3,500
C=A - B Depreciable base $ 19,500
D Life [in years] 5
E=C/D Annual SLM depreciation $ 3,900
F=E/C SLM Rate 20.00%
G=F x 2 DDB Rate 40.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
2018 $ 23,000 40.00% $ 9,200 $ 9,200 $ 13,800
2019 $ 13,800 40.00% $ 5,520 $ 14,720 $ 8,280